# LaMP-QA: A Benchmark for Personalized Long-form Question Answering

### v1

@@ Abstract
Personalization is essential for question answering systems that are user-centric. Despite its importance, personalization in answer generation has been relatively underexplored. This is mainly due to lack of resources for training and evaluating personalized question answering systems. We address this gap by introducing LaMP-QA—a benchmark designed for evaluating personalized long-form answer generation. The benchmark covers questions from three major categories: (1) Arts & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture, encompassing over 45 subcategories in total. To assess the quality and potential impact of the LaMP-QA benchmark for personalized question answering, we conduct comprehensive human and automatic evaluations, to compare multiple evaluation strategies for evaluating generated personalized responses and measure their alignment with human preferences. Furthermore, we benchmark a number of non-personalized and personalized approaches based on open-source and proprietary large language models (LLMs). Our results show that incorporating the personalized context provided leads to performance improvements of up to 39%. The benchmark is publicly released to support future research in this area.
@@
Alireza Salemi  and Hamed Zamani Center for Intelligent Information Retrieval (CIIR) University of Massachusetts Amherst {asalemi,zamani}@cs.umass.edu Data & Benchmark: https://hf.co/datasets/alireza7/LaMP-QA Codes & GitHub: https://github.com/LaMP-Benchmark/LaMP-QA
@@ 1 Introduction
Personalization plays a key role in many applications such as search Xue et al. (2009), recommendation Naumov et al. (2019), and text generation Salemi et al. (2024b, 2025b); Xu et al. (2025), as it contributes to improving user satisfaction and fostering trust in the system. In the context of information seeking, personalization is valuable as it enables systems to generate responses that are tailored to the intent, background, and preferences of the user, producing more accurate and user-specific responses. Research on personalized information seeking has predominantly focused on personalized retrieval Kasela et al. (2024); Guo et al. (2021); Eugene et al. (2013), while the generation aspect of the problem remains underexplored. This gap has become increasingly important with the advent of LLMs that interact with users through natural language, making personalized text generation a critical area of study. Recently, researchers developed LaMP Salemi et al. (2024b) and LongLaMP Kumar et al. (2024) benchmarks for personalized text generation, however, they exclusively focus on content generation tasks, such as email generation based on user’s writing style, overlooking information seeking tasks. This leaves a critical gap in evaluating how well LLMs can generate tailored responses to users’ information needs.
@@
To assess the benchmark’s quality for personalization, we run a series of experiments. First, we assess the quality of automatically extracted rubric aspects—used as evaluation rubrics—from the user’s question narratives. Human annotators assign an average rating of 4.9 out of 5 on the quality of the extracted aspects, demonstrating the high quality of the aspect extraction process. We compare our proposed personalized, aspect-based evaluation method against two alternative strategies: pairwise comparison and aspect-free scoring (both also based on the question narrative). Our method achieves the highest alignment with human judgment, validating its effectiveness as an evaluation approach. Moreover, an essential characteristic of a high-quality personalized QA dataset is that the included user profile should yield better performance when used with the corresponding user’s query compared to either (1) no personalization, or (2) using a profile from another user. Our results confirm this: using the target user’s profile improves performance by up to 39% over the non-personalized setting, and by up to 62% compared to using a mismatched profile. These findings show that both the profiles and evaluation rubrics are user-specific and effectively support personalization. To establish baselines, we evaluate a range of open-source and proprietary LLMs—including Gemma 2 Gemma-Team (2024), Qwen 2.5 Qwen et al. (2025), and GPT-4o OpenAI (2024)—in both personalized (using RAG) and non-personalized settings on the LaMP-QA benchmark. All data and code is publicly released to encourage further research in personalized question answering. [^1] Code and data available at: https://github.com/LaMP-Benchmark/LaMP-QA
@@ 3.1 Data Collection
Method Arts & Lifestyle & Personal Society & Entertainment Development Culture train validation test train validation test train validation test #Questions (users) 9349 801 767 7370 892 989 7614 810 1074 #Evaluation Aspects $2.7\pm 0.9$ $4.7\pm 1.2$ $4.6\pm 1.2$ $3.1\pm 1.0$ $5.1\pm 1.1$ $5.1\pm 1.2$ $2.9\pm 0.9$ $4.8\pm 1.1$ $4.8\pm 1.0$ Profile Size $106.7\pm 127.3$ $129.0\pm 183.7$ $159.1\pm 203.0$ $116.6\pm 162.0$ $98.2\pm 198.6$ $111.6\pm 220.3$ $141.3\pm 194.7$ $110.5\pm 210.6$ $115.8\pm 203.6$ Question Length $13.0\pm 2.9$ $10.6\pm 4.0$ $10.0\pm 3.8$ $13.6\pm 3.3$ $11.3\pm 4.4$ $11.6\pm 4.6$ $14.2\pm 3.6$ $12.1\pm 4.9$ $12.9\pm 5.4$ Narrative Length $113.1\pm 98.2$ $166.1\pm 167.6$ $144.7\pm 146.0$ $132.2\pm 104.1$ $159.2\pm 138.5$ $169.4\pm 145.2$ $144.6\pm 117.9$ $161.6\pm 158.2$ $167.9\pm 143.4$
@@
We use the SE-PQA train, validation, and test sets, applying the filtering steps to retain only questions that are suitable for personalized evaluation. This ensures the dataset includes questions that benefit from personalization and contain aspects reflecting the user’s specific information needs. For each remaining post from user $u$, we treat the post’s question as the input $x_{u}$ and extract the relevant aspects, denoted as $E_{x_{u}}$, from the question narrative $r_{x_{u}}$. To construct the user profile $P_{u}$, we gather previous posts from the same user, capturing their information-seeking behavior. For a more fine-grained evaluation, we categorize the filtered questions into three categories: (1) Art & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture. Each of the main categories consists of subcategories, totaling over 45 subcategories. The distribution of subcategories is shown in Figure 3 in Appendix A. The dataset statistics are provided in Table 1. Two examples from our dataset, which includes the question, question narrative, and the extracted key aspects relevant to the user, is presented in Figures 6 and 7 in Appendix A.
@@ 5.2 Main Findings
The performance of the baselines on the LaMP-QA benchmark is reported in Table 2 for the test set and in Table 3 in Appendix F for the validation set. The results in Table 2 demonstrate that Plan-RAG Personalization significantly outperforms all baselines across all evaluation categories. This highlights the utility of the training data provided by the LaMP-QA benchmark for training an effective planner model that can infer the key information users expect in response to their questions. Furthermore, the results show that all personalized baselines leveraging the asker’s profile to tailor the LLM’s response outperform the non-personalized model in nearly all cases. This indicates that incorporating user-specific context enhances the relevance and quality of the generated responses for each user, underscoring the importance of personalization in question answering. Finally, we observe that the best performance is achieved when GPT-4o-mini is used as the backbone LLM for generation. Notably, the highest performance occurs when Plan-RAG Personalization is applied to this model, highlighting the effectiveness of the planning step in understanding user preferences and incorporating them into the generated response. This demonstrates that even with a strong underlying LLM, explicitly modeling user intent through planning provides substantial gains in personalized question answering.
@@
One might ask about alternative methods for evaluating generated personalized responses. We compare three approaches, each aiming to assess how well a response aligns with the user’s information needs. First, inspired by prior work Kocmi and Federmann (2023); Liu et al. (2023), we use a direct scoring approach in which, given the question, the question narrative representing the user’s stated information need, and the generated response, the LLM assigns a score between 0 and 1 using the prompt shown in Figure 8 in Appendix C. Second, we evaluate responses in a pairwise setup, where the LLM is provided with the question, the user’s stated information need, and two candidate responses, and is asked to choose the better one using the prompt in Figure 9. Lastly, we use our proposed evaluation method in Section 3.2, which scores responses based on how well they address the individual aspects extracted from the user’s information needs. The implementation details for all evaluation approaches are provided in Appendix C.
@@ 7 Conclusion
This paper introduced the LaMP-QA benchmark, specifically designed to evaluate personalized question answering with LLMs. The benchmark spans three broad domains: Arts & Entertainment, Lifestyle & Personal Development, and Society & Culture. We conducted a comprehensive analysis to assess the benchmark’s effectiveness in facilitating the evaluation of personalization in question answering. We investigate multiple evaluation strategies and find that aspect-based evaluation achieves the highest alignment with human judgment. We evaluate standard RAG baselines and introduce novel planning-based for generating personalized responses. Experimental results demonstrate that leveraging personalized context leads to substantial improvements in response quality.
@@ Appendix A Filtering Prompts & Examples from the LaMP-QA Benchmark
To illustrate the structure of questions, question narrative, and rubric aspects in the LaMP-QA benchmark, Figure 6 and Figure 7 present two representative examples from the dataset. These examples demonstrate how the benchmark captures user-specific information needs and the corresponding personalized evaluation criteria.
@@ Appendix C Evaluation Metric Details
5:     $s^{t}_{e_{i}}=\pi(x_{u},r_{u},e_{i},\hat{y}_{x_{u}})$triangleright Scoring output based on aspect using prompt in Figure 10
6:     $s_{e_{i}}=\frac{s^{t}_{e_{i}}}{2}$triangleright Normalizing the aspect score for aspect by division by 2
7:     $s^{t}_{\hat{y}_{x_{u}}}=s^{t}_{\hat{y}_{x_{u}}}+s_{e_{i}}$triangleright Score accumulation for averaging
@@
9:$s_{\hat{y}_{x_{u}}}=\frac{s^{t}_{\hat{y}_{x_{u}}}}{|E_{x_{u}}|}$triangleright Averaging the output score using division by the number of aspects
10:return $s_{\hat{y}_{x_{u}}}$triangleright Returning score for output for user $u$
@@ Appendix G AI Assistance Usage
In this paper, ChatGPT [^10] https://chat.openai.com/ has been used as a writing assistant. In more detail, an initial paragraph is given to the ChatGPT and asked to paraphrase the given text. Further edits were also applied to the generated text and then used.

### v2

@@ Abstract
Personalization is essential for question answering systems that are user-centric. Despite its importance, personalization in answer generation has been relatively underexplored. This is mainly due to lack of resources for training and evaluating personalized question answering systems. We address this gap by introducing LaMP-QA—a benchmark designed for evaluating personalized long-form answer generation. The benchmark covers questions from three major categories: (1) Arts & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture, encompassing over 45 subcategories in total. To assess the quality and potential impact of the LaMP-QA benchmark for personalized question answering, we conduct comprehensive human and automatic evaluations, to compare multiple evaluation strategies for evaluating generated personalized responses and measure their alignment with human preferences. Furthermore, we benchmark a number of non-personalized and personalized approaches based on open-source and proprietary large language models. Our results show that incorporating the personalized context provided leads to up to 39% performance improvements. The benchmark is publicly released to support future research in this area.
@@
Alireza Salemi  and Hamed Zamani Center for Intelligent Information Retrieval University of Massachusetts Amherst {asalemi,zamani}@cs.umass.edu Data: https://hf.co/datasets/alireza7/LaMP-QA Code: https://github.com/LaMP-Benchmark/LaMP-QA
@@ 1 Introduction
Personalization plays a key role in many applications such as search Xue et al. (2009), recommendation Naumov et al. (2019), and generation Salemi et al. (2024b, 2025b); Xu et al. (2025), as it contributes to improving user satisfaction and trust in the system. For information seeking, personalization is valuable as it enables systems to generate responses that are tailored to the intent, background, and preferences of the user, producing more accurate and user-specific responses. Research on personalized information seeking has predominantly focused on personalized retrieval Kasela et al. (2024); Guo et al. (2021); Eugene et al. (2013), while the generation aspect of the problem remains underexplored. This gap has become increasingly important with the advent of large language models (LLMs) that interact with users through natural language, making personalized text generation a critical area of study. Recently, researchers developed LaMP Salemi et al. (2024b) and LongLaMP Kumar et al. (2024) benchmarks for personalized text generation, however, they exclusively focus on content generation tasks, such as email generation based on user’s writing style, overlooking information seeking tasks. This leaves a critical gap in evaluating how well LLMs can generate tailored responses to users’ information needs.
@@
To assess the benchmark’s quality for personalization, we run a series of experiments. First, we assess the quality of automatically extracted rubric aspects—used as evaluation rubrics—from the user’s question narratives. Human annotators assign an average rating of 4.9 out of 5 on the quality of the extracted aspects, demonstrating the high quality of the aspect extraction process. We compare our proposed personalized, aspect-based evaluation method against two alternative strategies: pairwise comparison and aspect-free scoring (both also based on the question narrative). Our method achieves the highest alignment with human judgment, validating its effectiveness as an evaluation approach. Moreover, an essential characteristic of a high-quality personalized QA dataset is that the included user profile should yield better performance when used with the corresponding user’s query compared to either (1) no personalization, or (2) using a profile from another user. Our results confirm this: using the target user’s profile improves performance by up to 39% over the non-personalized setting, and by up to 62% compared to using a mismatched profile. These findings show that both the profiles and evaluation rubrics are user-specific and effectively support personalization. To establish baselines, we evaluate a range of open-source and proprietary LLMs—including Gemma 2 Gemma-Team (2024), Qwen 2.5 Qwen et al. (2025), and GPT-4o OpenAI (2024)—in both personalized (using RAG) and non-personalized settings on the LaMP-QA benchmark. All data and code is publicly released to encourage further research in personalized question answering. [^1] Code is available at: https://github.com/LaMP-Benchmark/LaMP-QA and the data is available at: https://hf.co/datasets/alireza7/LaMP-QA
@@ 3.1 Data Collection
Method Arts & Lifestyle & Personal Society & Entertainment Development Culture train validation test train validation test train validation test #Questions (users) 9349 801 767 7370 892 989 7614 810 1074 #Evaluation Aspects $2.7\pm 0.9$ $4.7\pm 1.2$ $4.6\pm 1.2$ $3.1\pm 1.0$ $5.1\pm 1.1$ $5.1\pm 1.2$ $2.9\pm 0.9$ $4.8\pm 1.1$ $4.8\pm 1.0$ Profile Size $106.7\pm 127.3$ $129.0\pm 183.7$ $59.1\pm 203.0 11$$16.6\pm 162.0$ $98.2\pm 198.6$ $111.6\pm 220.3$ $41.3\pm 194.7 11$$0.5\pm 210.6 11$$15.8\pm 203.6$ Question Length $13.0\pm 2.9$ $10.6\pm 4.0$ $10.0\pm 3.8$ $3.6\pm 3.3 11$$3\pm 4.4 11.$$1.6\pm 4.6$ $14.2\pm 3.6$ $12.1\pm 4.9$ $12.9\pm 5.4$ Narrative Length $113.1\pm 98.2$ $166.1\pm 167.6$ $144.7\pm 146.0$ $132.2\pm 104.1$ $159.2\pm 138.5$ $169.4\pm 145.2$ $144.6\pm 117.9$ $161.6\pm 158.2$ $167.9\pm 143.4$
@@
We use the SE-PQA train, validation, and test sets, applying the filtering steps to retain only questions that are suitable for personalized evaluation. This ensures the dataset includes questions that benefit from personalization and contain aspects reflecting the user’s specific information needs. For each remaining post from user $u$, we treat the post’s question as the input $x_{u}$ and extract the relevant aspects, denoted as $E_{x_{u}}$, from the question narrative $r_{x_{u}}$. To construct the user profile $P_{u}$, we gather previous posts from the same user, capturing their information-seeking behavior. For a more fine-grained evaluation, we categorize the filtered questions into three categories: (1) Art & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture. Each of the main categories consists of subcategories, totaling over 45 subcategories. The distribution of subcategories is shown in Figure 5 in Appendix A. The dataset statistics are provided in Table 1. Two examples from our dataset, which includes the question, question narrative, and the extracted key aspects relevant to the user, is presented in Figures 8 and 9 in Appendix A.
@@ 5.2 Main Findings
The performance of the baselines on the LaMP-QA benchmark is reported in Table 2 for the test set and in Table 3 in Appendix F for the validation set. The results in Table 2 demonstrate that Plan-RAG Personalization significantly outperforms all baselines across all evaluation categories. This highlights the utility of the training data provided by the LaMP-QA benchmark for training an effective planner model that can infer the key information users expect in response to their questions. Furthermore, the results show that all personalized baselines leveraging the asker’s profile to tailor the LLM’s response outperform the non-personalized model in nearly all cases. This indicates that incorporating user-specific context enhances the relevance and quality of the generated responses for each user, underscoring the importance of personalization in question answering.
Moreover, we observe that the best performance is achieved when GPT-4o-mini is used as the backbone LLM for generation. Notably, the highest performance occurs when Plan-RAG Personalization is applied to this model, highlighting the effectiveness of the planning step in understanding user preferences and incorporating them into the generated response. This demonstrates that even with a strong underlying LLM, explicitly modeling user intent through planning provides substantial gains in personalized question answering.
To further assess the comparative quality of generated responses, we conducted a human evaluation between the two strongest baselines, PlanPers and RAG-Personalization, both with Qwen 2.5 Instruct as the backbone. We randomly sampled 100 outputs from each system and asked two human annotators to evaluate them. For each instance, annotators were provided with the question, its narrative, and the personalized rubrics, and were asked to select the better response or indicate a tie. The results, shown in Figure 1, indicate that PlanPers was preferred in 35% of cases, while RAG-Personalization was preferred in 26% of cases. The remaining instances were judged as ties. These findings suggest that PlanPers produces responses that are more consistently aligned with the question narrative and personalized rubrics from a human perspective.
Figure 1: Results of pairwise human evaluation between RAG-Personalization and PlanPers with Qwen 2.5 Instruct 7B as the backbone LLM. Each slice reports the winning percentage of the corresponding method.
@@
Headroom Analysis for Improving Planning in PlanPers.
To evaluate the potential upper bound of plan quality in PlanPers, we conduct an experiment comparing our trained planner with an oracle setting. In the oracle condition, instead of generating plans using the learned planner, we directly use gold-standard plans derived from the rubric aspects associated with each user. These gold plans are then provided to the model to generate responses. The results of this comparison, shown in Figure 3 for Qwen 2.5 Instruct, indicate that access to gold plans substantially improves performance, yielding gains of 155%–208% on different tasks. While the trained planner already achieves significant improvements over baseline methods (Table 2), this gap highlights considerable headroom for further enhancing the planner’s ability to generate high-quality and user-specific plans.
Figure 3: Comparison between trained planner and utilizing gold plans on the performance of PlanPers (Qwen 2.5 Instruct 7B) on the test set of LaMP-QA.
@@
One might ask about alternative methods for evaluating generated responses. We compare three approaches, each aiming to assess how well a response aligns with the user’s information needs. First, inspired by prior work Kocmi and Federmann (2023); Liu et al. (2023), we use a direct scoring approach in which, given the question, the question narrative representing the user’s stated information need, and the generated response, the LLM assigns a score between 0 and 1 using the prompt shown in Figure 10 in Appendix C. Second, we evaluate responses in a pairwise setup, where the LLM is provided with the question, the user’s stated information need, and two candidate responses, and is asked to choose the better one using the prompt in Figure 11. Lastly, we use our proposed evaluation method in Section 3.2, which scores responses based on how well they address the individual aspects extracted from the user’s information needs. The implementation details for all evaluation approaches are provided in Appendix C.
@@ 7 Conclusion
This paper introduced the LaMP-QA benchmark, specifically designed to evaluate personalized question answering with LLMs. The benchmark spans three broad domains: Arts & Entertainment, Lifestyle & Personal Development, and Society & Culture. We conducted a comprehensive analysis to assess the benchmark’s effectiveness in facilitating the evaluation of personalization in question answering. We investigate multiple evaluation strategies and find that aspect-based evaluation achieves the highest alignment with human judgment. We evaluate standard RAG baselines and introduce novel planning-based methods for generating personalized responses that align more with the user’s information needs. Experimental results demonstrate that leveraging personalized context alongside planning-based personalized response generation leads to substantial improvements in response quality and personalization.
Acknowledgment
@@ Appendix A Filtering Prompts & Examples from the LaMP-QA Benchmark
To illustrate the structure of questions, question narrative, and rubric aspects in the LaMP-QA benchmark, Figures 8 and 9 present two representative examples from the dataset. These examples demonstrate how the benchmark captures user-specific information needs and the corresponding personalized evaluation criteria.
@@ Appendix C Evaluation Metric Details
5:  $s^{t}_{e_{i}}=\pi(x_{u},r_{u},e_{i},\hat{y}_{x_{u}})$ $\triangleright$ Scoring output based on aspect using prompt in Figure 12
6:  $s_{e_{i}}=\frac{s^{t}_{e_{i}}}{2}$ $\triangleright$ Normalizing the aspect score for aspect by division by 2
7:  $s^{t}_{\hat{y}_{x_{u}}}=s^{t}_{\hat{y}_{x_{u}}}+s_{e_{i}}$ $\triangleright$ Score accumulation for averaging
@@
[…]
10:return $s_{\hat{y}_{x_{u}}}$ $\triangleright$ Returning score for output for user $u$
@@ Appendix G AI Assistance Usage
We used ChatGPT [^10] https://chat.openai.com/ as a writing assistant. Specifically, initial drafts of certain paragraphs were paraphrased using ChatGPT, after which manual revisions were applied before inclusion in the paper.